Jesse goes to the store. He buys a magazine for $8. Then he sells it for $10 and buys it back again for $11. He finally sells it for $12. What was his profit?
A) $-1
B) $0
C)$2
D$3

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Jesse goes to the store. He buys a magazine for $8. Then he sells it for $10 and buys it back again for $11. He finally sells it for $12. What was his profit?
================
He paid $19
He received $22.

Answer by ikleyn(52781)

.

His profit was  (10-8)  +  (12-11) = 2 + 1 = 3 dollars.